dy/dx=sin(x+y)+cos(x+y)

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear upendra

dy/dx=sin(x+y)+cos(x+y)

let   x+y=u

1+dy/dx=du/dy

so differential become

du/dy -1 =sinu  +cosu

du/dy=sinu +cosu +1

du/(sinu +cosu +1)   =dy

du/(sinu +cosu +1) =∫dy

now this a ordinary intigral u can easly solve it


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin